OpenStudy (anonymous):

Simplify the expression and explain each step. 12+3)2y-3

OpenStudy (skullpatrol):

$$\huge 12+3)2y-3$$ This is not a meaningful expression.

OpenStudy (anonymous):

How would I find y?

OpenStudy (skullpatrol):

Look at the expression, are you missing a (

OpenStudy (anonymous):

yes, sorry

OpenStudy (anonymous):

12+3(2y-3

OpenStudy (mew55):

is it 12+3(2y-3)?

OpenStudy (anonymous):

So i'm thinking just get rid of everything that you can (12+3) and then distribute

OpenStudy (mew55):

hmmm. letz c. 12+3(2y-3)=0 15(2y-3)=0 30y-45=0 30y=45 y=1.5

OpenStudy (mew55):

is it right?

OpenStudy (anonymous):

well it says to simplify not solve for y 12+3(2y-3) distribute the 3 12+6y-9 combine terms 6y+3 and it's simplified

OpenStudy (anonymous):

it don't have what it equals so we can't solve for y